a 14 litre solution is 25% antifreeze and 75% water, how much antifreeze should i add to make the solution 30% antifreeze?
Answer by ptaylor(2198)

Let x=amount of antifreeze that needs to be added
Now we know that the amount of water before the mixture takes place is equal to the amount of water after the mixture takes place, so:
0.75*14=0.70(14+x)
10.5=9.8+0.70x subtract 9.8 from each side
10.5-9.8=0.70x
0.7=0.70x
x=1 gal --------------amount of pure antifreeze needed
Another way: The amount of pure antifreeze also remains the same before and after the mixture takes place, so:
0.25*14+x=0.30(14+x)
3.5+x=4.2+0.30x
x-0.30x=4.2-3.5
0.70x=0.7
x=1 gal
Hope this helps---ptaylor